OpenStudy (anonymous):

Choose one of the 50 states at random. a.) what is the probability that it begins with M?????????

OpenStudy (anonymous):

12 states start with a vowel.

OpenStudy (anonymous):

12/50 = 6/25 = 0.24

OpenStudy (anonymous):

you have a 12 out of 50 = 12/50 chance, since there are 12 states starting with vowels. I was just simplifying the fraction to 6/25 (dividing numerator and denominator by 2).

OpenStudy (anonymous):

this the answer i got when i asked it

OpenStudy (anonymous):

i thot that was part b) i dont get wat u do for part a)

OpenStudy (anonymous):

well it also depends how many of the 50 states altogether do begin with a letter M it self.

OpenStudy (anonymous):

yea i was thinkin that..but how wud the equation be done to solve for this question

OpenStudy (anonymous):

oh yea im sorry thats part b

OpenStudy (anonymous):

there are 8state that start with M

OpenStudy (anonymous):

so take 8/50= 0.16 and that is 16%

OpenStudy (anonymous):

ohh lol dats it..ok thank you